OpenStudy (anonymous):

Given sin0=12/37 and 90degrees < 0 < 180degrees , find the exact value of cos 0/2 .

OpenStudy (anonymous):

\[\theta\]

OpenStudy (luigi0210):

Draw a triangle

OpenStudy (anonymous):



OpenStudy (anonymous):

Ok now what

OpenStudy (luigi0210):

Fill in the values you know and solve for the missing one.

OpenStudy (anonymous):

I don't understand trigonometry so i don't know where to put the angles at all

OpenStudy (luigi0210):

._. \[\LARGE sin \theta= \frac{opposite}{hypotenuse} \]

OpenStudy (anonymous):

Okay so now what *le stupid when it comes to trigonometry*

OpenStudy (luigi0210):

You know what, you don't really need the triangle.. just solve for the adjacent side with Pythagorean theorem.

OpenStudy (anonymous):

Ok one sec